Probability problem:
There are 5 red and 4 black balls in a box. If you pick out 3 balls without replacement, what is the probability of getiing at least one red ball?
I have tried so far: P(5/9) + P(4/9) - (3/9)= (6/9)≈ 2/3

Answer by longjonsilver(2297)

P(at least 1 red) is a lot of work. However, thankfully, we can think of it as 1 - P(3 black), which is a lot easier to compute.

P(BBB) = (4/9)*(3/8)*(2/7) = 1/21

so P(at least 1 red) = 1 - 1/21
P(at least 1 red) = 20/21

jon.